# Automatically Advancing LLM Expertise in Technology Judgment

### v3

@@ Abstract
Large language models (LLMs) are rapidly becoming core tools for science, engineering, and innovation. Their promise lies not just in remembering facts, but in putting knowledge to work. Despite their impressive ability to answer increasingly difficult questions, it remains unclear whether LLMs truly use their knowledge when confronted with new and challenging tasks. We address this question with a patent classification task that requires deep conceptual understanding: distinguishing objectively different but semantically similar patents. To evaluate this approach, we introduce a challenging new benchmark of 1.3 million post-2015 computer science patent pairs, characterized by dense technical jargon and strategically complex writing. We find that LLMs often fail our benchmark and struggle to distinguish among semantically similar patents. To probe this failure, we introduce a novel framework that decomposes model errors into two sources: missing and unused knowledge. Our approach asks models to generate clarifying questions to improve their understanding, and then compares three settings: raw performance, self-answered questions, and externally supplied answers. This decomposition reveals that LLMs often possess the relevant knowledge internally but fail to deploy it, while a smaller share of errors arises from genuine knowledge gaps. We then ask whether the ability of models to construct a task-specific database of questions and answers differs across models. We find that smaller models generate simpler, broadly transferable questions, while larger models propose more complex but less generalizable ones. This suggests new strategies for combining strengths across models. Taken together, our findings highlight a critical limitation of current LLMs and their evaluation: models often know more than they can use. By shifting evaluation from recall of static facts to application of dynamic knowledge, our approach provides a more informative lens on model capabilities and opens a path toward building systems that better support technological discovery and innovation.
@@ 1 Introduction
Large language models (LLMs) are increasingly understood not merely as passive text generators, but systems capable of reasoning about and engaging with complex realities. Recent studies have highlighted their potential for self-improvement—forming hypotheses, seeking clarification, and revising beliefs through iterative interactions (Huang et al., 2023; Qu et al., 2024; Song et al., 2024; Wang et al., 2023c, b). Yet, despite their impressive capabilities, LLMs often struggle with tasks that hinge on subtle conceptual distinctions (Asthana et al., 2024; Havlík, 2024; Pavlick, 2023). Much of their internal knowledge remains latent, poorly organized, and difficult to access (Pan et al., 2025; Hoang et al., 2023). This reflects a fundamental tension: while LLMs appear to “know everything,” their knowledge is stored in a heavily compressed form that impairs activation, recall, and information quality assessment (Hoang et al., ; Hoang et al., 2023; Jaiswal et al., 2024).
In this paper, we ask whether LLMs can enhance their own understanding by acting more like learners: generating and answering their own questions – often tailored to their specific uncertainties and usage, reflecting personalized internal gaps rather than universally missing facts – to identify and fill in the background knowledge required for task completion. This simple strategy of self-questioning, especially when paired with external retrieval, can automatically improve judgment, expose internal knowledge gaps, and activate reasoning processes that not only enhance the model’s own performance but also transfer effectively to larger models, where such reasoning often remains dormant under standard prompting.
A further challenge we highlight is the distinction between lay-in knowledge and working knowledge. In humans, once a fact is known, it is typically accessible for reasoning and problem solving: retrieval and use are closely aligned. In LLMs, however, the presence of knowledge in parameters does not guarantee its availability in practice. Prior work has probed this issue by designing tasks that isolate narrow domains (e.g., math word problems or factual recall) to demonstrate what knowledge LLMs are or are not using (Asthana et al., 2024; Havlík, 2024; Pavlick, 2023). By contrast, our framework is situated in a broad technological domain, where the challenge is not the absence of information but the difficulty of activating and applying it in context. This generalizable setup, built on a large and diverse patent corpus, goes beyond special-case demonstrations. It allows us to investigate systematically how LLMs bridge the gap between latent stored knowledge and the working knowledge required for fine-grained conceptual judgment.
Our first contribution is the proposal of a series of behavioral experiments to systematically evaluate self-questioning as a core mechanism for active learning and introspective growth in LLMs. We focus on the model’s understanding of technology concepts, operationalized through a challenging pairwise differentiation task. Instead of assessing understanding through conventional summarization or question-answer (QA) tasks, we present LLMs with pairs of semantically similar but conceptually distinct patent abstracts and ask whether they describe the same invention. A model that truly understands the material must go beyond surface-level similarity and detect subtle functional or mechanistic differences. To rigorously test this ability, we construct a new dataset of over 1.3 million close-but-distinct computer science patent pairs, granted post-2015 by the United States Patent and Trademark Office (USPTO), which is our second contribution. Patents present a uniquely demanding testbed: they are written in dense, legal-technical language, often strategically obfuscated, and their distinctions are adversarially verified by expert patent examiners. Our task requires fine-grained conceptual discrimination that goes well beyond pattern matching, skills, imitation, or memorization that could possibly bypass the existing benchmarks (McIntosh et al., 2024; Wu et al., 2024; Davis, 2023)  [^1] We have more detailed discussions in Appendices A and B.. We embed a structured self-questioning phase between initial exposure and final prediction. Given a pair of closely related patents, the model generates questions targeting both surface-level details and deeper conceptual understanding. It retrieves relevant scientific content to answer its own questions, combining external evidence with internal knowledge. These QA pairs, without verified correctness, are then fed back into the model to inform its final judgment on whether the patents describe the same invention. This setup allows us to assess whether and how self-generated inquiry and targeted retrieval enhance the model’s conceptual understanding (i.e., its accuracy).
Using our framework, we show that prompting LLMs with self-questioning significantly improves their performance on the patent differentiation task, enhancing judgment accuracy, confidence, and response consistency. We then explore three hypotheses to understand what drives this improvement. One possibility is that the questions themselves help the model parse the patent text more effectively by serving as informative cues. Another is that LLMs possess relevant internal knowledge but often fail to access it. A third hypothesis is that even when the model does retrieve relevant knowledge, its internal representations may compress and lose useful information, such that access to external science demonstrably improves performance. To test these possibilities, we design experiments to explore how LLMs’ internal and external knowledge are organized and transferred across model scales, especially when they are sparsely distributed like science, which is our third contribution. We compare the model’s accuracy when given only the questions, when using its questions and self-generated answers, and when using questions and answers derived from science. We find that prompting the model to ask itself questions helps it parse the patent text and improves performance. Feeding LLMs with their self-produced answers also helps, suggesting that LLMs often fail to retrieve relevant internal knowledge. We also find that self-generated answers are less effective than externally sourced ones, however, indicating that hidden compressed capacity remains insufficient for complex judgment tasks, where raw, external scientific knowledge in the form of QA proves beneficial to add richness. Interestingly, we find that small models often generate more useful, open-ended, and better-aligned questions for mid-sized models than larger models do, which highlights the importance of matching question complexity to model capacity, as well as a scalable prompting strategy where smaller models help structure the problem in a straightforward way for stronger models to reason through.
2 A Unique Dataset of Hard-to-Distinguish Patents
Understanding is a complex cognitive ability that goes beyond recalling facts or executing procedures. It involves grasping the deep relational structures that connect concepts, including causal links, analogies, and hierarchies. Drawing from cognitive science and philosophy, a central aspect of understanding is differentiation: recognizing what makes two concepts distinct. This ability is crucial for constructing models of conceptual relationships and is especially vital for causal reasoning, where detecting contrasts and similarities supports counterfactual thinking and explanatory modeling. [^2] Section A in the appendix offers a more detailed discussion of the link between understanding and pairwise differentiation.
Because direct questions like “what is X?” often elicit superficial or memorized responses, we need a more robust test of understanding. Instead, we propose a task that evaluates the model’s capacity to discriminate between concepts that are linguistically similar but semantically distinct. Our test centers on pairwise differentiation: the model is presented with two short texts and asked whether they describe the same underlying idea or two distinct ones. These texts are carefully selected to be semantically similar in language and structure — making the task deceptively difficult, and surface-level heuristics ineffective. The model must detect subtle differences in purpose, mechanism, and/or function, and reason about their implications.
We apply this task in the technology domain and focus specifically on patent abstracts. We construct a novel dataset consisting of 1,319,184 pairs of information and computer technology patent abstracts granted by the USPTO after 2015. In the Appendix C, we also explore another type of "merging" task: whether a model can recognize a rewritten description referring to the same invention. We find that the differentiation task requires genuine conceptual linking and functional comparison, making it a strong testbed for evaluating LLM understanding in the technology domain. Several features make technology patents particularly well-suited for testing fine-grained semantic distinctions:
@@ 2 A Unique Dataset of Hard-to-Distinguish Patents
(b) Patents are hard to distinguish
Figure 1: Panel (a): Model perplexity of LLaMA-3.1-8B-Instruct on paper and patents abstracts. Panel (b): Histogram of all 1,319,184 patents, showing the distribution of each patent’s maximum cosine similarity to another (distinct) patent.
First, patents often describe components of larger systems and therefore exist in dense clusters of near-duplicate-documents that are nearly identical in structure and language but differ in one or two meaningful ways, such as a specific constraint, use case, or technical improvement (Trippe, 2015). These subtle distinctions have been rigorously evaluated by human judges (we will develop this in the third point) and represent precisely the type of contrasts that test whether an LLM can move beyond surface-level similarity and recognize functional difference. We illustrate this pattern in Figure 1(b), which shows the distribution of the most similar—but non-identical—patent pairs based on textual embeddings from (Ghosh et al., 2024), pretrained on patent-citation relations. The distribution is heavily concentrated at high similarity scores, suggesting that most patents have at least one semantically similar counterpart.
Second, patent documents are often strategically obfuscated, written to disclose the minimum required for legal protection while revealing as little as possible about the underlying invention. As a result, their language tends to be intentionally dense, formal, and legally contorted. Readability studies show that patent specifications—particularly the claims sections—require significantly higher grade-level comprehension than comparably technical scientific prose, due to their use of dense jargon and legalistic drafting norms (Kong et al., 2023; Ouellette, 2012). This linguistic complexity is further illustrated in Figure 1(a), which compares the model perplexity of LLaMA-3.1-8B-Instruct on 4,000 randomly selected computer science paper abstracts from ArXiv and 4,000 matched patent abstracts from the same years. Perplexity, while not a direct measure of understanding, reflects the model’s inability to predict the next word in a sequence. Consistently higher perplexity on patent abstracts suggests that they contain highly technical and domain-specific language that poses a greater challenge for LLMs to parse and comprehend. An example of paired patent abstracts and a related paper abstract can be seen in Appendix D.
Third, and most importantly, even though patents may seem similar, we know they are not. Each granted patent in our dataset has undergone an extensive expert evaluation process conducted by patent examiners at the USPTO. This review assesses whether the application satisfies the legal requirements of novelty, non-obviousness, and utility. During this process, examiners conduct detailed prior art searches, compare the proposed invention to existing patents and publications, and determine whether the claimed invention constitutes a meaningful and non-trivial contribution over what already exists. This institutional validation is crucial for our task: it provides a strong, domain-informed guarantee that two different granted patents indeed describe conceptually distinct inventions.
@@ 3 Evaluating LLMs’ Understanding of the Technology Domain
We now describe how we operationalize the pairwise differentiation task. We begin by sampling pairs of patents that are difficult to distinguish, identified by embedding all patent abstracts using PATCER (Ghosh et al., 2024), a contrastive learning model trained on patent citation relationships. For each patent, we select its most similar—but non-identical—counterpart based on cosine similarity in the embedding space. These pairs are then presented to the LLM, which is asked whether the two abstracts describe the same invention. [^3] See Appendix E for full prompts. For each prompt, the model returns two outputs: a binary label indicating “yes” (1) or “no” (0), and a confidence score on a scale from 0 to 10.
@@
To construct our main measure of understanding, we aggregate predictions using a confidence-weighted voting scheme applied over three independent generations of the same prompt. If the total confidence associated with label “1” exceeds that of label “0,” the final label is set to 1; otherwise, it is 0. This aggregation method has been shown to represent the model’s actual certainty, accelerate convergence toward the model’s stable judgment, and improve self-consistency relative to simple majority voting (Geng et al., 2024). In addition to providing more reliable final decisions, confidence scores offer a fine-grained signal that helps distinguish between borderline and clear-cut cases. This is particularly important in our setting, where many patent pairs are deceptively similar—confidence levels can help identify when the model is genuinely uncertain (Geng et al., 2024). Alongside final labels, we also track changes in confidence and measure self-consistency, defined as the proportion of times all three responses yield the same answer. Lower consistency indicates greater uncertainty in model judgment.
@@
Figure 2: Ask LLMs "Do you think they are the same patent/different patents?", given that they are actually different. All models are Llama 3.1 instruction-tuning versions.
Figure 2 illustrates how models of different sizes perform on the differentiation task within the technology domain. The first panel shows that smaller models, such as LLaMA-8B, struggle to distinguish between semantically similar but distinct patents, classifying over 70% of patent pairs as the same. In contrast, the larger LLaMA-405B model labels only around 10% of these pairs as identical, indicating a stronger ability to detect subtle distinctions. The second and third panels show that the smaller model is also less confident and consistent in its responses (across blue bars). The 8b model’s confidence and self-consistency of wrong answers (blue) is way higher than those of the right answer (orange), again suggesting higher uncertainty and less stable internal representations compared to the larger model. The fourth panel evaluates this phenomenon directly by measuring whether the model gives consistent answers across two question forms. For each patent pair, we ask both “Do you think these patents describe the same invention?” and “Do you think they describe different inventions?” A model with a stable internal representation of the concepts should give logically consistent answers across these two formulations, measured by joint-correct rate, the proportion of pairs where the model gives both correct answers. Deviations from this indicate sensitivity to surface-level wording. The results again show that smaller models are less consistent and more affected by such superficial cues in the phrasing, reinforcing the idea that they lack a robust conceptual understanding of the task.
In the appendix F.1, we show that across all LLaMa models, responses shift systematically depending on how the question is framed. All models are more likely to judge the patents as different when asked explicitly whether they are different, even though the content of the question is logically equivalent. We also report additional results for other model families with varying sizes, including Qwen, Mixtral, and GPT-4o-mini. Again, we find that larger models perform better on the differentiation task. Interestingly, we observe consistent response biases in some models: Mixtral models tend to answer “yes” regardless of the question’s phrasing, while Qwen models tend to answer “no.” Based on these findings, we select the LLaMA-3.1-8B Instruct model as the primary focus for our subsequent experiments and use the question “Do you think they are the same patent?” as the default formulation for the pairwise differentiation task.
4 Self-questioning LLMs for Improved Technology Understanding

### v4

# The Missing vs. Unused Knowledge Hypothesis for Language Model Bottlenecks in Patent Understanding

@@ Abstract
While large language models (LLMs) excel at factual recall, the real challenge lies in knowledge application. A gap persists between their ability to answer complex questions and their effectiveness in performing tasks that require that knowledge. We investigate this gap using a patent classification problem that requires deep conceptual understanding to distinguish semantically similar but objectively different patents written in dense, strategic technical language. We find that LLMs often struggle with this distinction. To diagnose the source of these failures, we introduce a framework that decomposes model errors into two categories: missing knowledge and unused knowledge. Our method prompts models to generate clarifying questions and compares three settings – raw performance, self-answered questions that activate internal knowledge, and externally provided answers that supply missing knowledge (if any). We show that most errors stem from failures to deploy existing knowledge rather than from true knowledge gaps. We also examine how models differ in constructing task-specific question–answer databases. Smaller models tend to generate simpler questions that they, and other models, can retrieve and use effectively, whereas larger models produce more complex questions that are less effective, suggesting complementary strengths across model scales. Together, our findings highlight that shifting evaluation from static fact recall to dynamic knowledge application offers a more informative view of model capabilities.
The Missing vs. Unused Knowledge Hypothesis for Language Model Bottlenecks in Patent Understanding
@@ 1 Introduction
Large Language Models (LLMs) are increasingly viewed not merely as passive text generators, but as systems that can reason about and engage with complex realities. Yet despite their impressive capabilities, LLMs often struggle with tasks that hinge on subtle conceptual distinctions (Asthana et al., 2024; Havlík, 2024; Pavlick, 2023). We hypothesize that this difficulty reflects not a lack of knowledge, but limits in how knowledge is represented and accessed: much of what LLMs encode remains latent, poorly organized, and difficult to retrieve in knowledge-intensive settings (Pan et al., 2025; Hoang et al., 2023). This creates a fundamental tension: although LLMs can appear to “know everything,” their knowledge is stored in a highly compressed form that can hinder activation, recall, and information-quality assessment (Hoang et al., ; Hoang et al., 2023; Jaiswal et al., 2024). More broadly, the gap between the knowledge an LLM contains and what it reliably deploys at inference time is central to evaluating its competence. In humans, knowledge that is learned is typically accessible for reasoning and problem-solving; in LLMs, by contrast, having information in parameters does not guarantee its availability when needed. Prior work has examined this gap using narrowly scoped probes, such as math word problems or factual recall, to test what models do or do not use (Asthana et al., 2024; Havlík, 2024; Pavlick, 2023), but such tasks often miss the holistic, integrated nature of conceptual understanding.
We refer to these two forms of knowledge as lay in knowledge, knowledge the model can produce when directly prompted, and working knowledge, the knowledge it actually brings to bear in the course of solving a task.
To quantify the contribution of each type of knowledge to task performance, we design a conceptual understanding test that requires deploying knowledge dynamically rather than merely stating it. Understanding is a core cognitive capacity underlying most knowledge-intensive behavior: it goes beyond recalling facts or executing procedures, and instead involves grasping structural links and hierarchies that support transfer and generalization. Without some levels of understanding, both humans and LLMs may succeed on familiar patterns yet fail under distributional shift, novel recombinations of known components, or adversarially phrased problems. We therefore treat understanding not as an optional “nice-to-have,” but as a prerequisite for turning stored information into usable working knowledge. Our testbed is patent understanding, where success depends on fine-grained access to mechanisms, purposes, and constraints. [^1] Section A in the appendix offers a more detailed discussion of the link between understanding and pairwise differentiation.
Drawing from cognitive science and philosophy, we posit that a key, observable facet of understanding is differentiation – the ability to recognize what makes two concepts distinct. [^2] We provide a more detailed discussion in Appendices A and B. Accordingly, we evaluate in this paper models’ capacity to discriminate between complex technological concepts that are linguistically similar yet functionally distinct. In our task, a model is presented with two texts (patent descriptions) and asked whether they describe the same underlying idea or two distinct ones. These texts are carefully selected to be semantically similar in language and structure, making the task deceptively difficult and surface-level heuristics ineffective. To succeed, the model must detect subtle differences in aspects such as purpose and mechanism. We choose patents as our primary testbed because they present a uniquely demanding environment; they are written in dense, legal-technical language, often strategically obfuscated, and their distinctions are adversarially verified by human patent examiners. This task requires fine-grained conceptual discrimination that goes well beyond the pattern matching (in strategically confusing texts), imitation, or memorization that often bypasses existing benchmarks (McIntosh et al., 2024; Wu et al., 2024; Davis, 2023).
Our results show that while frontier models display some competence, weaker models consistently miss key technological nuances. Building on this, we quantify whether failures on knowledge-intensive understanding tasks stem from missing knowledge or from having “lay-in” knowledge that models cannot reliably use.
To measure the gap between working knowledge and lay-in knowledge, we prompt models to generate clarification questions to improve understanding about patents and compare baseline performance to two conditions: (i) questions and answers generated by the model itself, and (ii) questions and answers generated using external scientific information. We find that understanding failures are driven primarily by poor retrieval of lay-in knowledge: lay-in knowledge is substantially broader than working knowledge and, for our task, is nearly complete, containing roughly as much useful information as externally retrieved knowledge. Limited understanding is mainly a failure to access relevant knowledge at the right moment, not a failure to store it.
We then examine whether models differ in their ability to construct task-specific “databases” of usable cues. Smaller models tend to generate simpler, more broadly transferable cues that aid retrieval and downstream use, whereas larger models generate richer but more brittle and less generalizable cues. Together, these results point to a complementary strategy across scales: small models propose the scaffold, and large models do the heavy reasoning.
Contributions
(i) We introduce a scalable task and dataset to probe LLMs’ ability to deploy domain knowledge by operationalizing a core cognitive criterion of understanding - differentiation. We introduce a novel large-scale dataset of hard-to-separate USPTO patents (1.3M+ information/computer and 177K+ biomedical patents granted post-2015), where surface similarity is systematically misleading and success requires technical judgment. (ii) We introduce a diagnostic framework that separates missing from unused, "lay-in", knowledge. Applying this framework reveals a deployment bottleneck: models often have relevant knowledge but fail to activate it.
2 Hard-to-Distinguish Patents
The dataset consists of 1,319,184 information and computer technology and 177,903 biomedical patents. The designed task is less likely to be impacted by pretraining as shown in Appendix G. Several features make technology patents particularly well-suited for our test:
First, patents describe components of larger systems and thus occur in dense clusters of near-duplicates that are highly similar in structure and language but differ in a small number of substantively meaningful dimensions (e.g., constraints, use cases, or technical improvements) (Ouellette, 2011). These fine-grained distinctions - rigorously evaluated by human judges (see the third point) - directly test whether LLMs can move beyond surface similarity to recognize functional differences. We illustrate this pattern in Figure 1(b), which shows the distribution of similarities between one patent and its most similar – but non-identical – counterpart based on the textual embedding from (Ghosh et al., 2024), pretrained on patent-citation relations. The distribution is heavily concentrated at high similarity scores, suggesting that most patents have at least one semantically similar counterpart. We collect these similar but different patents and test whether LLMs can tell the nuances. In the Appendix C, we also explore another type of "merging" task: whether a model can recognize a rewritten description referring to the same invention. We find that the differentiation task requires genuine conceptual linking and functional comparison, making it a strong testbed for evaluating LLM understanding in the technology domain.
@@ 2 Hard-to-Distinguish Patents
(b) Patents and counterparts are hard to distinguish
Figure 1: Panel (a): Model perplexity of LLaMA-3.1-8B-Instruct on paper texts and patents descriptions. The texts of patents are much more confusing to LLMs compared to scientific papers. Panel (b): Histogram of the distribution of each patent’s maximum cosine similarity to another (distinct) patent.
Second, patent documents are often strategically obfuscated, disclosing only what is legally required while minimizing substantive revelation. Consequently, their language is deliberately dense, formal, and legalistic. Readability studies show that patent texts demand substantially higher comprehension than comparably technical scientific prose due to jargon-heavy and legally constrained drafting conventions (Kong et al., 2023; Ouellette, 2011). This linguistic complexity is further illustrated in Figure 1(a), which compares the model perplexity of LLaMA-3.1-8B-Instruct on 4,000 randomly selected computer science paper texts from ArXiv and 4,000 matched patent descriptions from the same years. Perplexity, while not a direct measure of understanding, reflects the model’s inability to predict the next word in a sequence. Consistently higher perplexity on patents suggests that they contain highly technical and domain-specific language that poses a greater challenge for LLMs to parse and comprehend. An example of paired patents and a related paper can be seen in Appendix D.
Third, and most importantly, despite surface similarity, granted patents are known to be conceptually distinct because each has undergone extensive expert review by USPTO examiners. This process evaluates novelty, non-obviousness, and utility through detailed prior-art searches and comparative analysis. As a result, institutional validation provides a strong, domain-informed, human-expert-evaluated guarantee that different granted patents represent non-trivial conceptual distinctions, even when they appear nearly identical.
@@ 3 Evaluating LLMs’ Understanding of the Technology Domain
Hard-to-distinguish patent pairs (two patents show high text similarity but with different granted IDs) are presented to the LLM, which is asked whether the two texts describe the same invention. [^3] See Appendix E for full prompts. For each prompt, the model returns two outputs: a binary label indicating “yes” (1) or “no” (0), and a confidence score on a scale from 0 to 10. Note that the expert-evaluated ground truth is always "they are different patents".
@@
To construct our main measure of understanding, we aggregate LLM’s output using a confidence-weighted voting scheme applied over three independent generations of the same prompt. If the total confidence associated with label “1” exceeds that of label “0,” the final label is set to 1; otherwise, it is 0. This aggregation method has been shown to represent the model’s actual certainty, accelerate convergence toward the model’s stable judgment, and improve self-consistency relative to simple majority voting (Geng et al., 2024; Taubenfeld et al., 2025). In addition to providing more reliable final decisions, confidence scores offer a fine-grained signal that helps distinguish between borderline and clear-cut cases. This is particularly important in our setting, where many patent pairs are deceptively similar—confidence levels can help identify when the model is genuinely uncertain (Geng et al., 2024). The result of confidence-weighted voting correlates with majority voting without any confidence signals, and the correlation is 0.95. High confidence also well correlates with high judge accuracy, with a correlation of 0.92. Results using the majority voting can be found in Appendix F. We also track changes in confidence. Throughout the experiments we use 5000 randomly sampled pairs to approximate the result on the full dataset as they are statistically sufficient. We find results on 1000, 5000, and 10000 randomly sampled pairs show a high correlation of over 0.98.
@@
Figure 2: Blue bars always represent the answer "yes", orange bars always represent "no". Top row: Ask LLMs "Do you think they are different patents?", given that they are actually different. Blue: "yes" (they are different). Orange: "no" (they are the same). Bottom row: "Do you think they are the same patent?", given that they are actually different. Blue: "yes" (they are the same). Orange: "no" (they are different). Llama models are instruction-tuning versions.
Looking at the top row of Figure 2, we evaluate LLM understanding by presenting two close-but-distinct patents and first asking <whether they are different?> Blue bars here always represent the answer "yes", and in this case, the proportion of responses "yes, they are different" (i.e., correct answer), along with their associated confidence, while the orange bars always denote the answer "no", and in this case, the responses "no, they are not different" (i.e., they are the same, incorrect answer). These results suggest that most models - with the exception of Qwen 2.5 7B - perform remarkably well; Mistral 7B, Llama 3.1 8B, and larger models like Llama 3.3 70B and GPT-4o mini all appear to identify the patents as distinct at a high rate (left panel, top row).
However, further investigation reveals that this seemingly high accuracy is superficial rather than a reflection of true understanding. We introduced a logically consistent inverse question in the bottom row of Figure 2: <Are these patents the same?> For a model to demonstrate a robust grasp of the nuances between these hard-to-distinguish patents, it should answer "yes" to the first prompt (are they different?) and "no" to the second (are they the same?). Interestingly, our results reveal significant response biases that undermine the top-row results. For instance, the Mistral model tends to answer "yes" regardless of the prompt’s framing, whereas Qwen 2.5 7B shows a persistent tendency to answer "no". Furthermore, we observe a paradoxical "overconfidence" effect, where small models (Qwen 2.5 7B, Mistral 7B, and Llama 3.1 8B) always exhibit higher confidence when they answer "yes", and for larger models like Llama 3.3 70b and GPT-4o-mini, wrong answers show a higher confidence than correct ones.
We also find GPT-4o mini demonstrates the strongest capability, followed by Llama 3.3 70B and then Llama 3.1 8B. This is evidenced by the fact that while all three models frequently identify that the patents are different ("yes" responses when asked if the patents are different) in the top row, GPT-4o mini maintains the lowest—and therefore most accurate—rate of "yes" responses when asked if the patents are the same in the bottom row. This suggests that as model scale and sophistication increase, so does the accuracy and the ability to maintain logical consistency when distinguishing between intricate technical documents.
4 Experimental setups
For failures in knowledge-intensive tasks, the most intuitive solution is to supplement the model’s knowledge, with the most direct method being Retrieval-Augmented Generation (RAG). Based on this premise, we designed a set of control experiments to analyze why models fail. Our motivation is straightforward: we allow the model to self-diagnose by asking, "What specific background knowledge would help you better distinguish between these two patents?"  [^4] See Appendix E for full prompts.. We then tested two distinct conditions: (1) Scientific RAG: given the set of generated questions, supplying external scientific knowledge to address missing information, if any. (2) Self-QA: given the same set of generated questions, LLMs answer their own questions, testing whether the model can bridge the gap using its own internal knowledge, indicating a utilization failure. We then compare the content of these two types of answers and their impact on improving the judgment to determine the major drivers of failures.
Specifically, we design two types of self-generated questions. The first, surface-level questions, target conceptual recall (e.g., What is X? What does Y do?), while the second encourages deeper understanding through questions about the integration of concepts, improvement, and the intent underlying invention. When the model is exposed to a pair of closely related patents, $P_{1}$ and $P_{2}$, it generates six questions $Q_{ij}$ for each, where $i\in{1,2}$ indexes the patent and $j\in{1,\dots,6}$ indexes the question. Three of these are surface-level questions, and three target deeper comprehension. Randomizing the order of questions does not impact the result much. The prompt for generating questions could be found in Appendix E.